# Large-Scale Multi-Robot Coverage Path Planning via Local Searchthanks: This work was supported by the NSERC under grant number RGPIN2020-06540 and a CFI JELF award. Code: https://github.com/reso1/LS-MCPP

### 17 Dec 2023 v1

@@ 1 Introduction
Figure 1: Graph-based CPP and MCPP: Gray squares, black circles, and black stars represent terrain graph vertices, decomposed graph vertices, and initial vertices of robots, respectively; Solid lines and dashed lines represent coverage paths and spanning edges, respectively. (a) The terrain to be covered where all terrain graph edges have uniform weights of $1$. (b) The single-robot coverage path generated by STC. (c)(d) Suboptimal and optimal $2$-robot coverage paths with makespans of $2$ and $1.5$, respectively.
@@
2 Related Work on MCPP
@@
3 Problem Formulation
In an MCPP instance, we are given a connected 2D grid terrain graph $G=(V_{g},E_{g})$ and its corresponding decomposed graph $D=(V_{d},E_{d})$, where each terrain vertex in $V_{g}$ is decomposed into four small adjacent vertices in $V_{d}$ as shown in Fig. 1-(a). For clarity, we let $\delta_{v}\in V_{g}$ denote the corresponding terrain vertex of a decomposed $v\in V_{d}$ and $\varepsilon=(\delta_{u},\delta_{v})\in E_{g}$ denote the edge connecting $\delta_{u}$ and $\delta_{v}$. Each $\varepsilon\in E_{g}$ has a weight of $w_{\varepsilon}$. Note that the definition of weighted edges is different from the literature that considers weighted vertices (Zheng et al. 2010), yet it is rather convenient to convert weighted edges to weighted vertices by evenly distributing each edge weight to its two endpoint vertices. In that case, we have $w_{\delta_{v}}=\sum_{\varepsilon\sim{\delta_{v}}}w_{\varepsilon}$ for each $\delta_{v}\in V_{g}$, where $\varepsilon\sim\delta_{v}$ represents that $\delta_{v}$ is an endpoint of $\varepsilon$. Each edge $e=(u,v)\in E_{d}$ of $D$ is thereby defined to have a weight $w_{e}=\frac{1}{4}\cdot\frac{1}{2}\cdot\left[\sum_{\varepsilon\sim\delta_{u}}w_{\varepsilon}+\sum_{\varepsilon\sim\delta_{u}}w_{\varepsilon}\right]$ by evenly splitting the weight $w_{\delta}$ of each $\delta\in V_{g}$ into $w_{\delta}/4$ for each of its four decomposed vertices (Tang and Ma 2023).
@@ 3 Problem Formulation
4 Extended-STC
In this section, we present the new Extended-STC (ESTC) paradigm to address coverage problems on incomplete terrain graphs. We define that $\delta_{v}\in V_{g}$ is complete if all its four decomposed vertices are present in $V_{d}$ and incomplete otherwise. A terrain graph $G$ is complete if all its vertices are complete and incomplete otherwise. STC works only for complete terrain graphs due to its reliance on circumnavigating spanning edges of $G$. As shown in Fig. 1-(c)(d), this restriction of STC can also prevent STC-based MCPP algorithms (Zheng et al. 2010; Tang and Ma 2023) from finding good solutions even on complete terrain graphs. In this paper, we focus on covering a connected but possibly incomplete decomposed graph, as covering an unconnected decomposed graph only additionally requires a robot in each of its connected components.
@@ 4 Extended-STC
$w_{\max}\cdot\frac{1}{2}\cdot\left({\textstyle\sum_{\varepsilon\sim\delta_{u}}w_{\varepsilon}+\sum_{\varepsilon\sim\delta_{u}}w_{\varepsilon}}\right)$ (1)
@@
5 Boundary Editing Operators
@@
6 LS-MCPP
In this section, we present LS-MCPP, a novel local search framework for effective MCPP solutions. LS-MCPP integrates the boundary editing operators described in Sec. 5, directly operating on the decomposed graph $D$ to explore good coverage paths by adjusting subgraphs $D_{i}$. We categorize $D_{i}$ as a light subgraph if its coverage path cost $c(\pi_{i})$ from ESTC is no larger than the average coverage path cost $\bar{c}$ of all subgraphs; otherwise, it is a heavy subgraph. LS-MCPP employs grow operators on light subgraphs and duplicate operators on heavy subgraphs to eliminate redundancy and exchange operators to balance path costs between subgraphs. The goal is to achieve a cost-equilibrium solution with minimal duplication. At a higher level, LS-MCPP employs a hierarchical sampling approach for efficient exploration of the constructed neighborhood in each iteration of its local search. It selects an operator pool using roulette wheel selection (Goldberg 1989) from three pools, each containing operators of the same type. Then, new heuristics are introduced to draw operators from the selected pool. LS-MCPP also calls a deduplication function periodically to exploit the current neighborhood and achieve a low-makespan solution.
@@ 6 LS-MCPP
Input: MCPP instance $(G,D,R)$, initial MCPP solution $\Pi$
@@ 3        Apply operator $o$ and update $\Pi$
6   $t\leftarrow\alpha\cdot t$ if $i\,\%\,S=0\vee\Delta\tau<0$ then
@@
8   Assign $\Pi$ to $\Pi^{*}$ if $\Pi$ is better than $\Pi^{*}$ Update each pool of $\mathcal{O}$
@@ 12            Remove $u,v$ from $D_{i}$ and update $\pi_{i}$
Forced Deduplication: A key consideration towards the success of LS-MCPP is to restrict duplication to only those necessarily needed over the neighborhood construction; for instance, a narrow path connecting two separate regions of the terrain in the flr-s instance (Tang and Ma 2023). For this purpose, forcedDeduplication() (line 1 of Alg. 1) aims to deduplicate all possible duplication in two folds. The first part (lines 1-1) iterates through each coverage path $\pi_{i}\in\Pi$ in a cost-decreasing order to remove any U-turn in it (lines 1-1). A U-turn is an edge $(u,v)\in\pi_{i}$ with $u,v\in V^{+}\cap V_{d,i}$, satisfying $\exists\,(p,q)\in\pi_{i}$ such that $(u,p),(v,q)\in\pi_{i}$. (so that $p\rightarrow u\rightarrow v\rightarrow q$ forms a “U-turn”.) The second part (lines 1-1) tries to recursively apply all valid deduplication operators in a descending order of the coverage path costs (line 1) and an ascending order of the heuristic values (line 1). Fig. 5 demonstrates forcedDeduplication() on a coverage path. It is worth noting that the first part can remove edges that do not constitute valid deduplication operators, and some valid deduplication operators can remove edges that are not U-turns (depending on the spanning tree generated by ESTC). The complementary relation between the two parts yields a better performance of the proposed LS-MCPP.
7 Empirical Evaluation
This section describes our experimental results on a Linux server with an Intel® Xeon® Gold 5218 CPU operating at 2.30 GHz and 300 GB memory limit. All LS-MCPP experiments were conducted across 12 repetitions with seeds spanning from 0 to 11.
[…]
@@ 7 Empirical Evaluation
Instances: We use MCPP instances from (Tang and Ma 2023), where their numbers of graph vertices, graph edges, and robots range from $46$ to $824$, $60$ to $1495$, and $4$ to $12$, respectively. As shown in Fig. 6 and Tab. 1, we design three additional larger instances whose terrain graphs are adopted from popular 2D pathfinding benchmarks (Sturtevant 2012). All the above instances have complete terrain graphs.
Baselines: We use MFC (Zheng et al. 2010), $*{}^{*}$ (Tang, Sun, and Zhang 2021), and MIP/MIP(H) (Tang and Ma 2023) as baseline algorithms that all use STC to transform a tree cover on $G$ to an MCPP solution. Both MFC and $*{}^{*}$ compute a suboptimal tree cover. MIP computes an optimal tree cover with a MIP model. MIP(H) computes a suboptimal tree cover with a reduced-size MIP model but is more efficient than MIP. We also design a new baseline VOR that first computes the Voronoi diagram (i.e., $k$ subgraphs) of $G$ based on the initial vertices of robots and then uses ESTC to generate coverage paths on the subgraphs.
Parameters: For all instances solved with LS-MCPP, the pool weight decay factor $\gamma$ is set to $0.01$, and the temperature decay factor $\alpha$ is set to $\exp{\frac{\log{0.2}}{M}}$ to decrease the temperature $t$ from $1$ to $0.2$. For instances from (Tang and Ma 2023), we set the max iteration $M$ to $33e$ 3 and the forced deduplication step $S$ to $21e$ 2, For the three additional larger instances in Tab. 1, we set $M$ to $41.5e$ 4 and $S$ to $25e$ 2.
7.1 Ablation Study
@@
Operator Sampling: We use a uniform random operator sampling method (Rand) to compare with the proposed heuristic-based operator sampling method (Heur) in Sec. 6. In Fig. 10, we compare the first and third columns, and the second and fourth columns. We can see that Heur is much better than Rand in flr-l; but it is only slightly better than Rand in trn1-l, possibly because trn1-l is a relatively simpler instance for LS-MCPP thus the difference between the two sampling methods is trivial.
@@
Forced Deduplication: In Fig. 10, we validate the function ForcedDeduplication (FD) of LS-MCPP in Sec. 6 by comparing the first and second columns, and the third and fourth columns. The periodic calling of FD makes the subgraphs return to states with only necessary duplications remaining, guiding LS-MCPP to better solutions.
7.2 Performance Comparison
@@
8 Conclusions and Future Work
We introduced LS-MCPP, a new local search framework that integrates the novel standalone ESTC paradigm and boundary editing operators to systematically explore MCPP solutions directly on decomposed graphs for the first time. Our experiments showcased its superiority over state-of-the-art baselines in reducing makespan and achieving complete coverage even for incomplete terrain graphs. The practicality of LS-MCPP was evident in large-scale scenarios, scaling efficiently in handling tens of thousands of vertices and dozens of robots in just minutes. Future work includes enhancing operator selection with machine learning techniques, speeding up LS-MCPP for larger-scale instances with parallelization techniques, and addressing potential inter-robot collisions in occluded environments with multi-agent pathfinding (Stern et al. 2019) techniques.

### 28 Feb 2024 v2

# Large-Scale Multi-Robot Coverage Path Planning via Local Searchthanks: Code: https://github.com/reso1/LS-MCPP

@@ Introduction
Figure 1: Graph-based CPP and MCPP: Gray squares, black circles, and black stars represent terrain graph vertices, decomposed graph vertices, and initial vertices of robots, respectively; Solid lines and dashed lines represent coverage paths and spanning edges, respectively. (a) Terrain graph with uniform edge weights. (b) The single-robot coverage path generated by STC. (c)(d) Suboptimal and optimal $2$-robot coverage paths with makespans $2$ and $1.5$, respectively.
@@
Related Work on MCPP
@@
Problem Formulation
In an MCPP instance, we are given a connected 2D grid terrain graph $G=(V_{g},E_{g})$ and its corresponding decomposed graph $D=(V_{d},E_{d})$, where each terrain vertex in $V_{g}$ is decomposed into four small adjacent vertices in $V_{d}$ as shown in Fig. 1-(a). For clarity, we let $\delta_{v}\in V_{g}$ denote the corresponding terrain vertex of a decomposed $v\in V_{d}$ and $\varepsilon=(\delta_{u},\delta_{v})\in E_{g}$ denote the edge connecting $\delta_{u}$ and $\delta_{v}$. Each $\varepsilon\in E_{g}$ has a weight of $w_{\varepsilon}$. Note that the definition of weighted edges is different from the literature that considers weighted vertices (Zheng et al. 2010), yet it is rather convenient to convert weighted edges to weighted vertices by evenly distributing each edge weight to its two endpoint vertices. In that case, we have $w_{\delta_{v}}=\sum_{\varepsilon\sim{\delta_{v}}}w_{\varepsilon}$ for each $\delta_{v}\in V_{g}$, where $\varepsilon\sim\delta_{v}$ represents that $\delta_{v}$ is an endpoint of $\varepsilon$. Each edge $e=(u,v)\in E_{d}$ of $D$ is thereby defined to have a weight $w_{e}=\frac{1}{4}\cdot\frac{1}{2}\cdot\left[\sum_{\varepsilon\sim\delta_{u}}w_{\varepsilon}+\sum_{\varepsilon\sim\delta_{v}}w_{\varepsilon}\right]$ by evenly splitting the weight $w_{\delta}$ of each $\delta\in V_{g}$ into $w_{\delta}/4$ for each of its four decomposed vertices (Tang and Ma 2023).
@@
Extended-STC
In this section, we present the new Extended-STC (ESTC) paradigm to address coverage problems on incomplete terrain graphs. We define that $\delta_{v}\in V_{g}$ is complete if all its four decomposed vertices are present in $V_{d}$ and incomplete otherwise. A terrain graph $G$ is complete if all its vertices are complete and incomplete otherwise. STC works only for complete terrain graphs due to its reliance on circumnavigating spanning edges of $G$. As in Fig. 1-(c)(d), this restriction of STC also prevents STC-based MCPP algorithms (Tang and Ma 2023) from finding good solutions even on complete terrain graphs. In this paper, we focus on covering a connected but possibly incomplete decomposed graph, as covering an unconnected decomposed graph only additionally requires a robot in each of its connected components.
@@
$w_{\max}\cdot\frac{1}{2}\cdot\left({\textstyle\sum_{\varepsilon\sim\delta_{u}}w_{\varepsilon}+\sum_{\varepsilon\sim\delta_{v}}w_{\varepsilon}}\right)$ (1)
@@
Boundary Editing Operators
@@
LS-MCPP
In this section, we present LS-MCPP, a novel local search framework for effective MCPP solutions. LS-MCPP integrates the boundary editing operators, directly operating on the decomposed graph $D$ to explore good coverage paths by adjusting subgraphs $D_{i}$. We categorize $D_{i}$ as a light subgraph if its coverage path cost $c(\pi_{i})$ from ESTC is no larger than the average coverage path cost $\bar{c}$ of all subgraphs; otherwise, it is a heavy subgraph. LS-MCPP employs grow operators on light subgraphs and deduplicate operators on heavy subgraphs to eliminate redundancy and exchange operators to balance path costs between subgraphs. The goal is to achieve a cost-equilibrium solution with minimal duplication. At a higher level, LS-MCPP employs a hierarchical sampling approach for efficient exploration of the constructed neighborhood in each iteration of its local search. It selects an operator pool using roulette wheel selection (Goldberg 1989) from three pools, each containing operators of the same type. Then, new heuristics are introduced to draw operators from the selected pool. LS-MCPP also calls a deduplication function periodically to exploit the current neighborhood and achieve a low-makespan solution.
@@
Input: MCPP instance $(G,D,R)$, initial solution $\Pi$
@@ 3        Apply operator $o$ and update $\Pi$
6   if $i\,\%\,S=0\vee\Delta\tau<0$ then
@@
8   Assign $\Pi$ to $\Pi^{*}$ if $\Pi$ is better than $\Pi^{*}$ Update each pool of $\mathcal{O}$ and assign $t$ to $\alpha\cdot t$
@@ 12            Remove $u,v$ from $D_{i}$ and update $\pi_{i}$
Forced Deduplication: A key consideration towards the success of LS-MCPP is to restrict duplication to only those necessarily needed over the neighborhood construction; for instance, a narrow path connecting two separate regions of the terrain in the flr-s instance (Tang and Ma 2023). For this purpose, forcedDeduplication() (line 1 of Alg. 1) aims to deduplicate all possible duplication in two folds. The first part (lines 1-1) iterates through each coverage path $\pi_{i}\in\Pi$ in a cost-decreasing order to remove any U-turn in it (lines 1-1). A U-turn is an edge $(u,v)\in\pi_{i}$ with $u,v\in V^{+}\cap V_{d,i}$, satisfying $\exists\,(p,q)\in\pi_{i}$ such that $(u,p),(v,q)\in\pi_{i}$. (so that $p\rightarrow u\rightarrow v\rightarrow q$ forms a “U-turn”.) The second part (lines 1-1) tries to recursively apply all valid deduplicate operators in a descending order of the coverage path costs (line 1) and an ascending order of the heuristic values (line 1). Fig. 5 demonstrates forcedDeduplication() on a coverage path. It is worth noting that the first part can remove edges that do not constitute valid deduplicate operators, and some valid deduplicate operators can remove edges that are not U-turns (depending on the spanning tree generated by ESTC). The complementary relation between the two parts yields a better performance of the proposed LS-MCPP.
Instance Grid spec. $\lvert V_{g}\rvert$ $\lvert E_{g}\rvert$ $k$ weighted
@@
Empirical Evaluation
This section describes our experimental results on an Intel® Xeon® Gold 5218 2.30 GHz Linux server with 300 GB memory. All LS-MCPP experiments were conducted across 12 repetitions with seeds spanning from 0 to 11. Instances: We use MCPP instances from (Tang and Ma 2023), where their numbers of graph vertices, graph edges, and robots range from $46$ to $824$, $60$ to $1495$, and $4$ to $12$, respectively. As shown in Fig. 6 and Tab. 1, we design three additional larger instances whose terrain graphs are adopted from popular 2D pathfinding benchmarks (Sturtevant 2012). All the above instances have complete terrain graphs. Baselines: We use MFC (Zheng et al. 2010), $*{}^{*}$ (Tang, Sun, and Zhang 2021), and MIP/MIP(H) (Tang and Ma 2023) as baseline algorithms that all use STC to transform a tree cover on $G$ to an MCPP solution. Both MFC and $*{}^{*}$ compute a suboptimal tree cover. MIP computes an optimal tree cover with a MIP model. MIP(H) computes a suboptimal tree cover with a reduced-size MIP model but is more efficient than MIP. We also design a new baseline VOR that first computes the Voronoi diagram (i.e., $k$ subgraphs) of $G$ based on the initial vertices of robots and then uses ESTC to generate coverage paths on the subgraphs. Parameters: For all instances solved with LS-MCPP, the pool weight decay factor $\gamma$ is set to $0.01$, and the temperature decay factor $\alpha$ is set to $\exp{\frac{\log{0.2}}{M}}$ to decrease the temperature $t$ from $1$ to $0.2$. For instances from (Tang and Ma 2023), we set the max iteration $M$ to $33e$ 3 and the forced deduplication step $S$ to $21e$ 2, For the three additional larger instances in Tab. 1, we set $M$ to $41.5e$ 4 and $S$ to $25e$ 2.
@@
Ablation Study
@@
Operator Sampling: We use a uniform random operator sampling method (Rand) to compare with the proposed heuristic-based operator sampling method (Heur). In Fig. 10, we compare the first and third columns, and the second and fourth columns. We can see that Heur is much better than Rand in flr-l; but it is only slightly better than Rand in trn1-l, possibly because trn1-l is a relatively simpler instance for LS-MCPP thus the difference between the two sampling methods is trivial.
@@
Forced Deduplication: In Fig. 10, we validate the function ForcedDeduplication (FD) of LS-MCPP by comparing the first and second columns, and the third and fourth columns. The periodic calling of FD makes the subgraphs return to states with only necessary duplications remaining, guiding LS-MCPP to better solutions.
Performance Comparison
@@
Conclusions and Future Work
We introduced LS-MCPP, a new local search framework that integrates the novel standalone ESTC paradigm and boundary editing operators to systematically explore MCPP solutions directly on decomposed graphs for the first time. Future work includes enhancing operator selection with machine learning techniques, speeding up LS-MCPP for larger-scale instances with parallelization techniques, and addressing potential inter-robot collisions in occluded environments with multi-agent pathfinding (Stern et al. 2019) techniques.
Acknowledgements
This work was supported by the NSERC under grant number RGPIN2020-06540 and a CFI JELF award.